### v1

@@ Abstract
Multilingual Large Language Models (LLMs) develop cross-lingual abilities despite being trained on limited parallel data. However, they often struggle to generate responses in the intended language, favoring high-resource languages such as English. In this work, we introduce CoCo-CoLa (Correct Concept - Correct Language), a novel metric to evaluate language adherence in multilingual LLMs. Using fine-tuning experiments on a closed-book QA task across seven languages, we analyze how training in one language affects others’ performance. Our findings reveal that multilingual models share task knowledge across languages but exhibit biases in the selection of output language. We identify language-specific layers, showing that final layers play a crucial role in determining output language. Accordingly, we propose a partial training strategy that selectively fine-tunes key layers, improving language adherence while significantly reducing computational cost. Our method achieves comparable or superior performance to full fine-tuning, particularly for low-resource languages, offering a more efficient multilingual adaptation. [^1] Our code is available at https://github.com/elnaz rahmati/CoCo-CoLa/
[Uncaptioned image]CoCo-CoLa: Evaluating Language Adherence in Multilingual LLMs
Elnaz Rahmatithanks: Equal contribution.                        Alireza S. Ziabari11footnotemark: 1                        Morteza Dehghani
University of Southern California {erahmati, salkhord, mdehghan}@usc.edu
@@ 1 Introduction
Multilingual LLMs are pre-trained on raw text from multiple languages, typically consisting of separate corpora for each language. Remarkably, despite this lack of explicit parallel data to facilitate cross-lingual associations, these models develop an implicit understanding of inter-language relations and cross-lingual word associations (Wen-Yi and Mimno, 2023). Instruction tuning further enhances their ability to follow prompts, and models trained on multilingual data often exhibit zero-shot cross-lingual transfer of instruction-following capabilities (Chirkova and Nikoulina, 2024). However, this generalization is uneven: while high-resource languages in pretraining benefit significantly from instruction tuning, lower-resource or unseen languages often struggle to follow instructions reliably, frequently exhibiting degraded performance or defaulting to generating output in a preferred language (Nguyen et al., 2024; Chirkova and Nikoulina, 2024). To address these issues, we investigate how multilingual LLMs learn the same task across different languages.
Figure 1: Evaluation of correctness and language adherence on French input. Soda level represents the CoCo-CoLa ratio. Higher level indicates better adherence.
@@
A crucial step toward addressing the limitations of multilingual LLMs is understanding how they internally process and encode multilingual knowledge. Interpretability research has traditionally focused on monolingual models, leveraging techniques such as representation probing (Orgad et al., 2024; Saphra and Lopez, 2019) and model patching (Ghandeharioun et al., 2024; García-Carrasco et al., 2024). These methods have been widely used to examine LLMs’ performance across tasks such as mathematics (Nikankin et al., 2024; Zhou et al., 2024), and general knowledge (Jiang et al., 2024; Burns et al., 2022; Singh et al., 2024; Golgoon et al., 2024; Chowdhury and Allan, 2024; Rai et al., 2024). Studies on model internals suggest that Multi-Layer Perceptrons (MLPs) retrieve task-relevant information, while attention layers refine and promote the correct response (Geva et al., 2021; Meng et al., 2022). Furthermore, knowledge is often identified in earlier layers and reinforced in later layers (Fan et al., 2024).
@@
However, these interpretability techniques have primarily been applied to monolingual models, which were initially dominant due to the early focus on English-language pertaining Touvron et al. (2023); Jiang et al. (2023); Team et al. (2024); Abdin et al. (2024). The rise of multilingual LLMs trained on diverse languages (Gao et al., 2024; Shaham et al., 2024; Soykan and Şahin, 2024), necessitates extending interpretability research beyond English. Multilingual LLMs present additional challenges: representations of different languages are intertwined within a shared space; cross-lingual alignment varies across languages; and shared tokens between languages impact their process. These complexities make it difficult to isolate language-specific knowledge, benchmark cross-lingual generalization, and interpret how multilingual LLMs acquire and apply linguistic information. Given the prevalence of mid- and low-resource languages, understanding these mechanisms is crucial not only for improving cross-lingual transfer but also for mitigating the “curse of multilinguality” — the performance degradation observed as the number of supported languages increases.
@@
Recent efforts have begun addressing these challenges by probing internal multilingual representations (Li et al., 2024), analyzing the emergence of cross-lingual transfer (Wang et al., 2024a), and studying token representation alignment on cross-lingual transfer (Gaschi et al., 2023). Furthermore, researchers attempt to separate the linguistic abilities from task-specific abilities by developing language- and task-specific adapters (Pfeiffer et al., 2020; Parovic et al., 2023), subnetworks (Choenni et al., 2023), or layers (Bandarkar et al., 2024). However, despite this progress, most prior works treat multilinguality as a monolithic phenomenon, focusing on general cross-lingual transfer or aggregating all languages into a single block of linguistic knowledge. Less attention has been given to understanding how LLMs process individual languages at a more granular level, particularly within the context of task learning.
@@
In this work, our goal is to identify both shared and distinct patterns in cross-lingual task acquisition, revealing how multilingual models internalize and apply linguistic knowledge (Section 3). We find that training on a task in one language improves performance in other languages. However, this benefit is not always directly observable due to an inherent model bias towards generating output in a preferred language, rather than strictly adhering to the input language (Section 4.1). To quantify this bias, we introduce CoCo-CoLa (Correct Concept, Correct Language), a novel metric designed to assess a model’s ability to generate responses in the intended input language, particularly for languages not included in supervised finetuning (SFT). Furthermore, we propose a partial training method that selectively fine-tunes specific model layers to enhance language adherence (Section 4.2). This approach enables more efficient multilingual adaptation, achieving comparable or even superior performance, especially for low-resource languages, without the need for full model retraining.
@@ 2 Related Work
Li et al. (2024) use probing techniques to analyze accuracy changes across layers in LLMs, showing that high-resource languages exhibit patterns similar to English, with accuracy increasing from lower to upper layers. However, this pattern is inconsistent for low-resource languages. Wang et al. (2024b) examine cross-lingual transfer by analyzing neuron overlap in different languages using checkpoints from BLOOM’s pre-training (Le Scao et al., 2023). Their study finds a strong correlation between neuron overlap and cross-lingual transfer, though neuron overlap does not increase monotonically during training, and patterns vary across model sizes. Similarly, Zhao et al. (2024a) investigate language-specific neurons and assess how masking these neurons affects both English and non-English language performance.
@@
Beyond studying multilingualism in LLMs, some research focuses on improving model performance across languages through representation alignment. Gaschi et al. (2023) align English and Arabic model representations using a bilingual dictionary before fine-tuning on a target task. Zhang et al. (2024) align English representations with other languages using question-translation data before instruction-tuning. Additionally, Salesky et al. (2023) introduce a pixel representation method to enhance alignment and improve translation quality.
@@
Another approach for cross-lingual transfer involves integrating adapters into the model. This technique is based on the assumption that task-solving knowledge can be separated from language knowledge. Pfeiffer et al. (2020) introduce MAD-X, a framework where language and task adapters are trained separately, with each block’s representations passing through a language adapter before a task adapter. Building on this, later works aim to refine adapter creation and composition methods. For instance, Parović et al. (2022) propose BAD-X, which replaces monolingual adapters with bilingual adapters, improving performance for low-resource languages. Zhao et al. (2024b) introduce AdaMergeX, where adapters for language-task pairs are trained independently and later combined through linear operations (addition and subtraction) to generate adapters for new language-task pairs.
@@
To enhance cross-lingual transfer without adding new parameters, some methods focus on identifying existing task- and language-specific parameters within the model. Choenni et al. (2023) fine-tune models for specific languages or tasks, extract the most affected neurons, and use the resulting subnetworks to enable multilingual task adaptation. Bandarkar et al. (2024) take a layer-wise approach in multiple steps: they train separate language- and task-expert models, analyze parameter changes to identify key layers for language and task learning, and use layer-swapping techniques to create a math expert in a new language. Consistent with Zhao et al. (2024a), their findings suggest that initial and final layers primarily encode language-related information, while middle layers are task-specific.
@@ 3 Preliminary Analysis
In the preliminary section of this paper, we first isolate language effects from task learning by choosing multi-lingual parallel QA data (Section 3.1), examining fine-tuning performance across multiple languages (Section 3.2), exploring how well LLMs generalize knowledge across languages (Section 3.3), and which model components are most affected during training (Section 3.4). Then, in Section 4.1, we introduce CoCo-CoLa metric to measure language adherence in multilingual LLMs followed by an efficient partial training method to increase the model adherence (Section 4.2).
@@ 3.1 Setup
To investigate how multilingual LLMs learn a new task in a monolingual setting, we train three different model sizes on a Closed-Book Question-Answering (CBQA) task. We include two sizes of the Llama-3.2 series (Dubey et al., 2024) to analyze the effect of model size on multilingual performance and behavior, given that these models are specifically optimized for multilingual dialogue. We also include Llama-3.1-8B as a point of comparison, as it, while not explicitly optimized for multilingualism, was trained on a small amount of multilingual data.
We select CBQA because it is inherently language-dependent and demonstrates a model’s ability to act as a knowledge base (Wang et al., 2021). To isolate the impact of language differences from the effects of learning a new task or acquiring new knowledge, we use the Mintaka CBQA dataset (Sen et al., 2022). Mintaka provides identical question-answer pairs in nine languages, allowing us to keep the question content consistent and thus isolate the influence of language itself. The dataset was originally created in English and later translated into Arabic, French, German, Hindi, Italian, Japanese, Portuguese, and Spanish.
@@ 3.2 SFT Performance
Our initial step is to assess the model’s ability to learn the task in each individual language, effectively measuring how learning difficulty varies across languages. To do this, we perform SFT for Llama-3 models on each language of the CBQA dataset for three epochs and generate answers for given questions. Next, we select the best model based on the validation loss. Further implementation details are provided in Section A.1.
Table 1 shows a comparison of accuracy between the pre-trained model and the best checkpoint of the language-specific SFT model across different languages. SFT significantly improves performance for all languages with relatively consistent accuracy levels, except for Hindi in all model sizes and Portuguese for Llama-3.1-8B, which exhibit notably lower accuracy. This discrepancy is likely due to undertraining. Among the SFT models, English achieves the highest accuracy in the 3B (53.09%) and 8B (50.98%) models, while Spanish performs best (41.71%) in the 1B model. The largest accuracy gains are observed in English (+38.06%) for the 8B model, German (+24.31%) for the 3B model, and Spanish (+35.58%) for the 1B model, indicating that these languages benefited the most from fine-tuning. The similar accuracy across languages indicates comparable knowledge acquisition.
However, two critical questions remain: (1) Do models share learned knowledge uniformly across languages, or do they correctly answer distinct subsets of questions depending on the language? (2) Are there specific parts of the model that are responsible for encoding language-specific information?
To address these questions, we first analyze the overlap in correct answers across languages using the Jaccard Index, followed by an investigation of parameter updates to determine whether certain components of the model specialize in handling linguistic differences.
Llama-1B Llama-3B Llama-8B
Language PLM SFT PLM SFT PLM SFT
English 13.27 38.44 32.85 53.09 12.92 50.98
French 11.30 40.27 22.90 43.80 18.53 50.85
German 7.16 40.34 23.79 48.10 11.04 44.35
Hindi 5.27 21.18 7.33 30.39 6.21 35.29
Italian 7.06 41.58 21.87 42.73 16.48 43.22
Portuguese 5.38 38.23 20.06 37.04 18.38 31.11
Spanish 6.13 41.71 22.01 45.69 16.60 45.46
Table 1: Performance of pre-trained (PLM) and fine-tuned (SFT) models across different languages.
@@ 3.3 Cross-lingual Task Knowledge
To further investigate the extent of cross-lingual task knowledge transfer within the model, we analyze the overlap in correct answers across languages. Specifically, we measure how consistently the model arrives at the same correct answers in different languages, providing insight into whether knowledge is shared across languages.
It is important to note that there is no overlap between the knowledge present in the training and evaluation data. This ensures that any correct answers during evaluation are derived from knowledge acquired during pretraining rather than memorization. Consequently, the model’s ability to generate correct responses across languages indicates that it has internalized the underlying task knowledge from the training data, rather than relying solely on language-specific cues. Let $L_{A}$ and $L_{B}$ represent two languages, and let $C_{L_{A}}$ denote the set of correct answers for $L_{A}$. To quantify the degree of shared task knowledge between languages, we compute the Jaccard Index, also known as Intersection over Union (IoU), between $C_{L_{A}}$ and $C_{L_{B}}$ (see Equation 1). The Jaccard Index is a natural choice for this analysis as it directly measures the proportion of overlapping correct answers relative to the total distinct answers across languages. This allows us to assess knowledge consistency and cross-lingual transfer within the model.
@@
The results, shown in Figure 2, indicate that on average approximately 60% of correctly answered questions are shared across languages for all models, suggesting a strong degree of shared knowledge among languages. However, Hindi exhibits significantly lower overlap with other languages in Llama-3.2 models, suggesting weaker generalization for this language. Interestingly, in Llama-3.1-8B, Hindi shows higher overlap compared to Llama-3.2 models, but Portuguese experiences a notable drop in overlap. Additionally, Llama-3.2-3B demonstrates a higher rate of shared knowledge compared to Llama-3.1-8B, despite both models achieving comparable accuracy across languages (see Table 1). This highlights the importance of multilingual optimization in enhancing cross-lingual transfer among languages.

### v2

@@ Abstract
Multilingual Large Language Models (LLMs) develop cross-lingual abilities despite being trained on limited parallel data. However, they often struggle to generate responses in the intended language, favoring high-resource languages such as English. In this work, we introduce CoCo-CoLa (Correct Concept - Correct Language), a novel metric to evaluate language adherence in multilingual LLMs. Using fine-tuning experiments on a closed-book QA task across seven languages, we analyze how training in one language affects others’ performance. Our findings reveal that multilingual models share task knowledge across languages but exhibit biases in the selection of output language. We identify language-specific layers, showing that final layers play a crucial role in determining output language. Accordingly, we propose a partial training strategy that selectively fine-tunes key layers, improving language adherence while reducing computational cost. Our method achieves comparable or superior performance to full fine-tuning, particularly for low-resource languages, offering a more efficient multilingual adaptation. [^1] Our code is available at https://github.com/elnaz rahmati/CoCo-CoLa/
@@ 1 Introduction
Multilingual LLMs are pre-trained on raw text from multiple languages, typically consisting of separate corpora for each language. Remarkably, despite this lack of explicit parallel data to facilitate cross-lingual associations, these models develop an implicit understanding of inter-language relations and cross-lingual word associations [1]. Instruction tuning further enhances their ability to follow prompts, and models trained on multilingual data often exhibit zero-shot cross-lingual transfer of instruction-following capabilities [2]. However, this generalization is uneven: while high-resource languages in pretraining benefit significantly from instruction tuning, lower-resource or unseen languages often struggle to follow instructions reliably, frequently exhibiting degraded performance or defaulting to generating output in a preferred language [3, 2]. To address these issues, we investigate how multilingual LLMs learn the same task across different languages.
Figure 1: Evaluation of correctness and language adherence on French input. The soda level visualizes the CoCo-CoLa ratio, with higher levels indicating stronger adherence to the input language. Our results show that partially fine-tuning the final layers of an English-tuned model on French achieves language adherence and accuracy comparable to a model fully fine-tuned on French.
@@
A crucial step toward addressing the limitations of multilingual LLMs is understanding how they internally process and encode multilingual knowledge. Interpretability research has traditionally focused on monolingual models, leveraging techniques such as representation probing [4, 5] and model patching [6, 7]. These methods have been widely used to examine LLMs’ performance across tasks such as mathematics [8, 9], and general knowledge [10, 11, 12, 13, 14, 15]. Studies on model internals suggest that Multi-Layer Perceptrons (MLPs) retrieve task-relevant information, while attention layers refine and promote the correct response [16, 17]. Furthermore, knowledge is often identified in earlier layers and reinforced in later layers [18].
@@
However, these interpretability techniques have primarily been applied to monolingual models, which were initially dominant due to the early focus on English-language pertaining [19, 20, 21, 22]. The rise of multilingual LLMs trained on diverse languages [23, 24, 25], necessitates extending interpretability research beyond English. Multilingual LLMs present additional challenges: representations of different languages are intertwined within a shared space; cross-lingual alignment varies across languages; and shared tokens between languages impact their process. These complexities make it difficult to isolate language-specific knowledge, benchmark cross-lingual generalization, and interpret how multilingual LLMs acquire and apply linguistic information. Given the prevalence of mid- and low-resource languages, understanding these mechanisms is crucial not only for improving cross-lingual transfer but also for mitigating the “curse of multilinguality” — the performance degradation observed as the number of supported languages increases.
@@
Recent efforts have begun addressing these challenges by probing internal multilingual representations [26], analyzing the emergence of cross-lingual transfer [27], and studying token representation alignment on cross-lingual transfer [28]. Furthermore, researchers attempt to separate the linguistic abilities from task abilities by developing language- and task-specific adapters [29, 30], subnetworks [31], or layers [32]. However, despite this progress, most prior works treat multilinguality as a monolithic phenomenon, focusing on general cross-lingual transfer or aggregating all languages into a single block of linguistic knowledge. Less attention has been given to understanding how LLMs process individual languages at a more granular level, particularly within the context of task learning.
@@
In this work, we focus on language adherence by first identifying both shared and distinct patterns in cross-lingual task acquisition, revealing how multilingual models internalize and apply linguistic knowledge (Section 3). We find that training on a task in one language improves performance in other languages. However, this benefit is not always directly observable due to an inherent model bias towards generating output in a preferred language, rather than strictly adhering to the input language (Section 4.1). To quantify this bias, we introduce CoCo-CoLa (Correct Concept, Correct Language), a novel metric designed to assess a model’s ability to generate responses in the intended input language, particularly for languages not included in supervised finetuning (SFT). Furthermore, we propose a partial training method that selectively fine-tunes specific model layers which reveals the relation between language adherence and model layers (Section 4.2). This approach enables more efficient language adaptation, achieving comparable or even superior performance compared to full model retraining, especially for low-resource languages. Finally, we show that the issue of language adherence can be addressed by finetuning only the final layers of LLMs on a small balanced multilingual data (Section 4.3).
@@ 2 Related Work
Li et al. [26] use probing techniques to analyze accuracy changes across layers in LLMs, showing that high-resource languages exhibit patterns similar to English, with accuracy increasing from lower to upper layers. However, this pattern is inconsistent for low-resource languages. Wang et al. [33] examine cross-lingual transfer by analyzing neuron overlap in different languages using checkpoints from BLOOM’s pre-training [34]. They find a strong correlation between neuron overlap and cross-lingual transfer, though neuron overlap does not increase monotonically during training, and patterns vary across model sizes. Similarly, Zhao et al. [35] investigate language-specific neurons and assess how these neurons affect both English and non-English language performance.
@@
Beyond studying multilingualism in LLMs, some research focuses on improving model performance across languages through representation alignment. Gaschi et al. [28] align English and Arabic model representations using a bilingual dictionary before fine-tuning on a target task. Zhang et al. [36] align English representations with other languages using question-translation data before instruction-tuning. Additionally, Salesky et al. [37] introduce a pixel representation method to enhance alignment and improve translation quality.
@@
Another approach for cross-lingual transfer involves integrating adapters into the model. This technique is based on the assumption that task-solving knowledge can be separated from language knowledge. Pfeiffer et al. [29] introduce MAD-X, a framework where language and task adapters are trained separately, with each block’s representations passing through a language adapter before a task adapter. Building on this, later works aim to refine adapter creation and composition methods. For instance, Parović et al. [38] propose BAD-X, which replaces monolingual adapters with bilingual adapters, improving performance for low-resource languages. Zhao et al. [39] introduce AdaMergeX, where adapters for language-task pairs are trained independently and later combined through linear operations (addition and subtraction) to generate adapters for new language-task pairs.
@@
To enhance cross-lingual transfer without adding new parameters, some methods focus on identifying existing task- and language-specific parameters within the model. Choenni et al. [31] fine-tune models for specific languages or tasks, extract the most affected neurons, and use the resulting subnetworks to enable multilingual task adaptation. Bandarkar et al. [32] take a layer-wise approach in multiple steps: they train separate language- and task-expert models, analyze parameter changes to identify key layers for language and task learning, and use layer-swapping techniques to create a math expert in a new language. Consistent with Zhao et al. [35], their findings suggest that initial and final layers primarily encode linguistic information, while middle layers are task-specific.
@@ 3 Preliminary Analysis
In the preliminary section of this paper, we first isolate language effects from task learning by choosing multi-lingual parallel QA data (Section 3.1), examine fine-tuning performance across multiple languages (Section 3.2), explore how well LLMs generalize knowledge across languages (Section 3.3), and which model components are most affected during training (Section 3.4). Then, in Section 4.1, we introduce CoCo-CoLa metric to measure language adherence in multilingual LLMs followed by an efficient partial training method to increase the model adherence (Section 4.2).
@@ 3.1 Setup
To investigate how multilingual LLMs learn a new task in a monolingual setting, we train four different models on a Closed-Book Question-Answering (CBQA) task. We include two sizes of the Llama-3.2 series [40] to analyze the effect of model size on multilingual performance and behavior, given that these models are specifically optimized for multilingual dialogue. We also include Llama-3.1-8B as a point of comparison, as it, while not explicitly optimized for multilingualism, was trained on a small multilingual corpus. To test generalizability to multilingualy balanced models, we include Gemma-3-4B [41], which was trained with UniMax [42] for addressing language imbalances.
We select CBQA because it is language-dependent and demonstrates a model’s ability to act as a knowledge base [43]. To isolate the impact of language differences from the effects of learning a new task or acquiring new knowledge, we use the Mintaka CBQA dataset [44]. Mintaka provides identical question-answer pairs in nine languages, allowing us to keep the question content consistent and thus isolate the influence of language itself. The dataset was originally created in English and later translated into Arabic, French, German, Hindi, Italian, Japanese, Portuguese, and Spanish.
@@ 3.2 SFT Performance
Table 1: Performance of pre-trained models (PLM), fine-tuned models (SFT), and their difference ($\Delta$ = SFT - PLM) on CBQA data across languages.
Llama-1B Llama-3B Llama-8B Gemma-4B
Language PLM SFT $\Delta$ PLM SFT $\Delta$ PLM SFT $\Delta$ PLM SFT $\Delta$
English 13.27 38.44 25.17 32.85 53.09 20.24 12.92 50.98 38.06 30.69 53.67 22.98
French 11.30 40.27 28.97 22.90 43.80 20.90 18.53 50.85 32.32 21.67 48.43 26.76
German 7.16 40.34 33.18 23.79 48.10 24.31 11.04 44.35 33.31 23.76 45.79 22.03
Hindi 5.27 21.18 15.91 7.33 30.39 23.06 6.21 35.29 29.08 8.84 43.96 35.12
Italian 7.06 41.58 34.52 21.87 42.73 20.86 16.48 43.22 26.74 20.44 50.05 29.61
Portuguese 5.38 38.23 32.85 20.06 37.04 16.98 18.38 31.11 12.73 19.96 44.16 24.20
Spanish 6.13 41.71 35.58 22.01 45.69 23.68 16.60 45.46 28.86 26.33 48.13 21.80
Our initial step is to assess the model’s ability to learn the task in each individual language, effectively measuring how learning difficulty varies across languages. To do this, we perform SFT for all models on each language of the CBQA dataset for three epochs and generate answers for given questions. Next, we select the best model based on the validation loss. Further implementation details are provided in Section A.3.
Table 1 shows a comparison of accuracy between the pre-trained model and the best checkpoint of the language-specific SFT model across different languages. SFT significantly improves performance for all languages with relatively consistent accuracy levels, except for Hindi in all Llama model sizes and Portuguese for Llama-8B, which exhibit notably lower accuracy. This discrepancy is likely due to undertraining. Among the SFT models, English achieves the highest accuracy in all models, except Llama-1B that performs best in Spanish. The largest accuracy gains are observed in English (+38.06%) for Llama-8B, German (+24.31%) for Llama-3B, Spanish (+35.58%) for Llama-1B, and Hindi (+35.12%) for Gemma-4B, indicating that these languages benefited the most from fine-tuning. The comparable accuracy across languages indicates comparable knowledge acquisition.
However, two critical questions remain: (1) Do models share learned knowledge uniformly across languages, or do they correctly answer distinct subsets of questions depending on the language? (2) Are there specific parts of the model that are responsible for encoding language-specific information? To address these questions, we first analyze the overlap in correct answers across languages using the Jaccard Index, followed by an investigation of parameter updates to determine whether certain components of the model specialize in handling linguistic differences.
@@ 3.3 Cross-lingual Task Knowledge
To further investigate the extent of cross-lingual task knowledge transfer within the model, we analyze the overlap in correct answers across languages. Specifically, we measure how consistently the model arrives at the same correct answers in different languages, providing insight into whether knowledge is shared across languages. It is important to note that there is no overlap between the knowledge present in the training and evaluation data. This ensures that any correct answers during evaluation are derived from knowledge acquired during pretraining rather than memorization. Consequently, the model’s ability to generate correct responses across languages indicates that it has internalized the underlying task knowledge from the training data, rather than relying solely on language-specific cues. Let $L_{A}$ and $L_{B}$ represent two languages, and let $C_{L_{A}}$ denote the set of correct answers for $L_{A}$. To quantify the degree of shared task knowledge between languages, we compute the Jaccard Index, also known as Intersection over Union (IoU), between $C_{L_{A}}$ and $C_{L_{B}}$ (see Equation 1). The Jaccard Index is a natural choice for this analysis as it directly measures the proportion of overlapping correct answers relative to the total distinct answers across languages. This allows us to assess knowledge consistency and cross-lingual transfer within the model.
@@
The results, shown in Figure 2, indicate that on average approximately 60% of correctly answered questions are shared across languages for all models, suggesting a strong degree of shared knowledge among languages. However, Hindi exhibits significantly lower overlap with other languages in Llama-3.2 models, suggesting weaker generalization for this language. Interestingly, in Llama-8B, Hindi shows higher overlap compared to Llama-3.2 models, but Portuguese experiences a notable drop in overlap. Additionally, Llama-3B demonstrates a higher rate of shared knowledge compared to Llama-8B, despite both models achieving comparable accuracy across languages (see Table 1). This highlights the importance of multilingual optimization in enhancing cross-lingual transfer among languages. For Gemma-4B, despite comparable accuracy across languages, the overall overlap is lower than that observed in the Llama models, indicating less cross-lingual knowledge sharing.